OpenStudy (anonymous):

Convert each rate using dimensional analysis. Round to the nearest tenth, if necessary. 45 mi/h = ___ ft/s a. 66 c. 26.4 b. 15.6 d. 19.2

OpenStudy (anonymous):

@blurbendy

OpenStudy (anonymous):

@itsonlycdeee

OpenStudy (johnweldon1993):

How many seconds are in a hour? 3600...so 1 unit conversion we'll be using is that... How many feet are in a mile? 5280..so that is the 2nd unit conversion we'll be using... \[\large \frac{ 45 \space \text{miles} }{ h } \times \frac{ 1 \space \text{Hour} }{ 3600 \space \text{Seconds} } \times \frac{ 5280 \space \text{Feet} }{ 1\space \text{Mile} }\] Can you solve that?